OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

You can just plug in x = 3 into those equations and see which one give you an answer of -2 in this case y = 2(3) - 8 = 6 -8 = -2